OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6x94 = 6x(?)+6x4 = (?)

ganeshie8 (ganeshie8):

looks like you're splitting \(94\) as \(90+4\) : \[\begin{align}6\times 94 &= 6\times (90+4) \\~\\&=6\times 90 + 6\times 4\end{align}\]

OpenStudy (anonymous):

which means

OpenStudy (anonymous):

can u tell me the answer

OpenStudy (anonymous):

the first question mark is 90 and the second is 564